### v2

# Hybrid Deep Learning Model to Estimate Cognitive Effort from fNIRS Signals in Educational Game Playing

Abstract.
This study estimates cognitive effort (CE) based on functional near- infrared spectroscopy (fNIRS) data and performance scores using a hybrid deep learning (DeepNet) model. The estimation of CE enables educators to modify material to enhance learning effectiveness and student engagement. In this study, we collected oxygenated hemoglobin $\Delta HbO$ using fNIRS during an educational quiz game. Participants (n=16) responded to 16 questions in a Unity-based educational game, each within a 30-second response time limit. The quiz was divided into two sessions, each session consisting of two task segments. Each segment had four questions. There was a short rest period between the two segments and a longer rest period after each session. We used machine learning and deep learning models to predict the performance score from the fNIRS signal, $\Delta HbO$ and compared traditional machine learning and DeepNet models to determine which approach provides better accuracy in predicting performance scores. The result shows that the GRU gives better performance with 67% than other models. After the prediction, we used the predicted score and the $\Delta HbO$ to observe cognitive effort by calculating relative neural efficiency and involvement in our test cases. Our result shows that even with moderate accuracy, the predicted CE closely follow the actual trends. This findings can be helpful in designing and improving learning environments and provide valuable insights into learning materials.
Deep learning, cognitive effort, relative neural efficiency, relative neural involvement, performance score, functional Near-Infrared Spectroscopy (fNIRS), brain signal, hemodynamic response, educational games
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: XXXXX; XX; isbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Applied computing Interactive learning environmentsccs: Applied computing Learning management systemsccs: Applied computing Computer-assisted instruction
Figure 1. The overview of the proposed system. (a) data collection: educational quiz gameplay was used to measure cognitive effort (CE) while recording fNIRS signals. (b) pre-processing: signals underwent pre-processing including filtering and $\Delta HbO$ extraction, and (c) models: various models including a hybrid deep learning model (CNN-GRU) was trained to predict performance scores from the signals(d) output: using the predicted scores and $\Delta HbO$, we estimated CE via Relative Neural Efficiency (RNE) and Involvement (RNI).
[…]
This teaser figure has four blocks showing the overall pipeline. In the leftmost picture (a), the data collection procedure has been showed. The image shows that the participant played an educational game, and we collected the score and oxygenated hemoglobin ($\Delta HbO$). We used an fNIRS device to collect $\Delta HbO$ and saved the score. Then, we combined all the data and created a dataset. The next part shows the steps of pre-processing (b). $\Delta HbO$ signals underwent pre-processing including filtering and $\Delta HbO$ extraction. (c) We showed that the dataset was sent to the models, and after that, we predicted the score (d). Then, based on the $\Delta HbO$ and score, we estimated the cognitive effort.
1. Introduction
Cognitive effort (CE) refers to the level of cognitive load an individual experiences while learning. This is associated with increased neural activity during problem-solving, decision-making, and learning. In recent years, understanding CE in education-based studies has increased (Getchell and Shewokis, 2023; Benerradi et al., 2019; Sharmin et al., 2024a). When students are learning new or complex tasks, measuring CE helps us understand how hard their brains work and how to improve the design of the educational content. Relative Neural Efficiency (RNE) and Relative Neural Involvement (RNI) are two metrics that can be used to explain CE with behavioral performance  (Getchell and Shewokis, 2023; Koiler et al., 2022; Reddy et al., 2022; Sharmin et al., 2024a). These metrics can be measured using hemodynamic response collected from the prefrontal cortex (PFC) using functional Near-Infrared Spectroscopy (fNIRS) and performance score during a task. The PFC is responsible for executive functions such as decision-making, memory retrieval, and cognitive flexibility, which are critical in educational contexts. RNE indicates how efficiently an individual processes information and performs a task with minimal CE. Higher RNE indicates low effort and high performance, and lower RNE suggests high effort and low performance. On the other hand, RNI measures motivation and engagement. High RNI indicates the user is more engaged and motivated during a task (Paas et al., 2003; Shewokis et al., 2015; Getchell and Shewokis, 2023; Koiler et al., 2022; Sharmin et al., 2024a) Thus, understanding RNE and RNI may help to optimize learning strategies, improve engagement, and enhance cognitive performance.
Traditional methods of evaluating student efficiency and involvement rely on test assessments, self-reports, or reaction times, which are often subjective and inconsistent (De La Cruz et al., 2022; Sharmin et al., 2024a, b). Also, a few studies measured students’ cognitive load, engagement, task difficulty, and efficiency based on only neuron activities(Pan et al., 2020; Sharmin and Barmaki, 2024; Sharmin et al., 2024b). Previously, different AI techniques have been applied in the field of brain signal analysis, such as cognitive workload estimation (Cao et al., 2022), pain level detection (Fernandez Rojas et al., 2019), brain-computer interfaces (BCI) controlling (Hamid et al., 2022), and mental health diagnosis (Kim et al., 2023b). These algorithms are capable of extracting complex patterns from neurophysiological data (Lu et al., 2020). A few recent studies used hemodynamic responses to understand learning efficiency along with subjective measures (Pan et al., 2020; Jeun et al., 2022; Oku and Sato, 2021). However, CE prediction using fNIRS and machine or DeepNet remains an open research challenge.
We derived RNE and RNI from z-scored performance and CE. We need both performance score and hemodynamic response, which offer a deeper lens into how hard participants work and how effectively they convert effort into learning. By predicting performance scores, we can make a connection between behavioral outcomes and neural measures. This could help improve instructional designs by identifying cognitive efficiency and engagement patterns.
By applying a Convolutional Neural Network- Gated Recurrent Unit (CNN-GRU) DeepNet model to estimate performance from fNIRS signals, we evaluate the possibility of predicting RNE and RNI, allowing the detection of cognitive fatigue, efficient learning, and recovery effects. To visualize, we divided the cartesian plot into four sub-parts. The High Effort + High Involvement (HE+HI) cognitive state is a critical learning phase observed during skill acquisition, especially in complex or multitasking environments. Individuals in the HE+HI quadrant exert substantial cognitive effort while remaining highly engaged with the task. This state often occurs in the early stages of training, where performance is still developing and mental workload is high due to unfamiliarity with the task or high task demands (Reddy et al., 2022). This quadrant should not be viewed negatively but rather as a productive struggle stage. In this phase, learners are cognitively investing in understanding the task structure, strategies, or switching between sub-tasks (Reddy et al., 2022). Over time, with sufficient practice, effective learners typically transition from HE+HI to LE+HI (Low Effort + High Involvement), showing improved neural efficiency while maintaining motivation (Reddy et al., 2022). Participants in the HE+HI quadrant often exhibit activation in task-relevant prefrontal areas, which supports task switching, attention, and working memory—further validating the cognitive demands being handled during this phase.
This study used hemodynamic responses collected by an fNIRS device to predict performance scores and measure CE by determining RNE and RNI in an educational game setting. Sixteen participants answered 16 questions in two quiz sessions, with fNIRS data recorded at a sampling rate of 10 Hz. We compare the effectiveness of various DeepNet models, such as CNN, LSTM, BiLSTM, and a hybrid CNN-GRU model for quiz score prediction. Then, we used these predicted scores and $\Delta HbO$ to estimate RNE and RNI.
Our research questions (RQs) are as follows:
  • RQ1: How does proposed CNN-GRU model compare to traditional ML approaches in predicting performance using brain signals?
  • RQ2 Can we determine the brain region most responsible for predicting cognitive effort using deep learning interpretability techniques on fNIRS data?
  • RQ3: Can a DeepNet model trained on fNIRS-derived features predict performance scores that meaningfully reflect cognitive effort through derived RNE and RNI metrics?
Contribution
While previous studies have applied machine learning and DeepNet models for cognitive workload estimation, they have primarily focused on classification tasks using standard methods such as n-back. These approaches often rely on subject-specific training and do not predict performance continuously, limiting their applicability in real-time learning environments. In contrast, our work introduces a novel dataset collected during an educational quiz game, capturing both fNIRS hemodynamic responses and performance scores. We predict continuous performance and estimate cognitive effort using relative neural efficiency (RNE) and involvement (RNI). Our hybrid CNN-GRU DeepNet model outperforms both traditional machine learning and standalone DL architectures. Furthermore, we demonstrate that GRU-derived features generalize well when used with XGBoost which suggests strong potential for cross-user adaptability.
The paper is organized as follows. section 2 reviews relevant literature on cognitive effort estimation, highlighting existing methods and identifying research gaps. section 3 outlines our experimental methodology, detailing participants, apparatus, study procedure, cognitive effort measurement methods, data acquisition procedures, and pre-processing techniques. In section 4, we present the proposed hybrid CNN-GRU DeepNet model architecture to evaluate the models. section 5 provides results comparing various machine learning and DeepNet models and analyzes predicted versus actual CE using RNE and RNI. section 6 discusses findings about the research questions, addressing practical implications, limitations, and possible improvements. Finally, section 7 summarizes the study’s findings and highlights its educational contributions.
2. Related Work
Hemodynamic responses using fNIRS have been used widely in cognitive workload assessment and improvement of controlling BCI applications. This is also used in neuro-degenerative disease studies such as Alzheimer’s, Attention-deficit/hyperactivity disorder (ADHD), autism, depression, etc. Machine learning (ML) and DeepNet (DL) have been applied in analyzing fNIRS data across various domains. Traditional ML models such as support vector machine (SVM) (Khan et al., 2021), k-nearest neighbors algorithm (KNN) (De La Cruz et al., 2022), linear discriminant analysis (LDA) (Grässler et al., 2021), etc., have been used for classification tasks, while DeepNet architectures like CNN and LSTM (Lu et al., 2020) have demonstrated superior accuracy in decoding brain activity. Researchers have also leveraged generative models (CNN-GANs) to address data scarcity issues and improve single-trial classification accuracy (Zhang et al., 2023). These advancements highlight the increasing role of ML/DL in fNIRS-based cognitive assessment that forms the foundation for applications in workload estimation and performance prediction.
fNIRS and Machine Learning in Brain Signal Analysis:
Deep and machine learning models are being applied to hemodynamic responses collected by fNIRS to control BCI applications by decoding speech (Cooney et al., 2022), motor imagery (Chiarelli et al., 2018; Ma et al., 2021), and finger tapping or movement (Khan et al., 2021; Ortega and Faisal, 2021; Jabbar et al., 2020). These algorithms are also used in health sector to classify pain (Fernandez Rojas et al., 2024; Eken et al., 2024), various neuro-degenerative diseases such as Alzheimer’s (Cicalese et al., 2020; Kim et al., 2023a), Parkinson’s (Quattrocelli et al., 2024; Guevara et al., 2024), ADHD (Shin et al., 2023; Zhao et al., 2021), depression (Li et al., 2025) etc. Detection of fatigue (Tanveer et al., 2019), cognitive load (Zhang et al., 2024), and cognitive tasks such as arithmetic task classification (Wickramaratne and Mahmud, 2021; Zhang et al., 2023) from fNIRS data using ML models gives satisfying results. DeepNet models, particularly CNN and long short-term memory (LSTM), have been effective in decoding brain activity from fNIRS signals with high classification accuracies. Wickramaratne and Mahmud proposed a CNN-based model for mental arithmetic classification and got 87.14% accuracy (Wickramaratne and Mahmud, 2021). Zhang et al. showed in their work that the CGAN-rIRN approach improves the single-trial accuracy (92.19%) for mental arithmetic and mental singing tasks in both the data augmentation and classifier (Zhang et al., 2023). Other models, such as the LSTM model for binary decision-making in BCI, achieved 83.89% accuracy (Erdoğan et al., 2024). Some studies have explored the integration of electroencephalogram (EEG) and fNIRS for improved performance in BCI applications (Cooney et al., 2022; Ortega and Faisal, 2021) by decoding speech (accuracy 87.18%) and moving hands (accuracy 54.4%) using DeepNet methods. EEG and fNIRS were also used in the health sector to classify Alzheimer’s disease (Cicalese et al., 2020), and mild cognitive impairment (Grässler et al., 2021) using LDA. These data were also used for working memory load classification using SVM (accuracy 100%) (Khan et al., 2021) and KNN (accuracy 72%) (De La Cruz et al., 2022).
Beyond these applications, fNIRS has been increasingly applied to predict cognitive workload, fatigue, and performance efficiency, as discussed in the next section.
fNIRS for Cognitive Workload, Fatigue, and Performance Prediction:
Beyond application in BCI and health sectors, researchers have increasingly explored how fNIRS-based models can predict cognitive workload and performance in real-world settings. fNIRS has been increasingly applied to predict cognitive workload and performance. Zhang et al. showed that DeepNet models can predict pilot workload in flight simulations using CNN-attention mechanisms with 88.69% accuracy (Zhang et al., 2024). Similarly, Grimaldi et al. proposed a CNN-LSTM model to predict perceptual load (Grimaldi et al., 2024a) and cognitive load (Grimaldi et al., 2024b) forecasting in piloting tasks. They stated that an LSTM-based approach demonstrated superior performance in short-term workload prediction (Grimaldi et al., 2024b). In another study, Gado et al. classified cognitive workload in NASA task-load (NASA-TLX) simulations using logistic regression, achieving 92.4% accuracy (Gado et al., 2023). Fatigue has also been investigated using fNIRS signals (Ma et al., 2025; Gado et al., 2023; Haroon et al., 2024; Varandas et al., 2022). For instance, fNIRS-based models were employed to classify cognitive fatigue during a digital learning task, achieving an accuracy of 70.91% ± 13.67% using Random Forest (RF)(Varandas et al., 2022). A study combining physiological signals with fNIRS for cognitive fatigue detection using RF reported 94.6% accuracy (Haroon et al., 2024). Ma et al. applied cnn-rESnET to detect fatigue and got 97.78% accuracy in their task (Ma et al., 2025). Most studies have focused on binary or categorical workload classification rather than continuous performance prediction.
In addition to workload and fatigue assessment, recent studies have explored the use of ML/DL for performance efficiency estimation, particularly in educational and learning environments. The following section discusses related works in this domain.
fNIRS in Performance and Efficiency Prediction:
While most studies focus on classifying cognitive workload, researchers have also begun investigating how ML/DL can predict performance efficiency and engagement based on fNIRS signals by monitoring neural activity in different learning and task-based environments.
Jeun et al. used logistic regression to predict task difficulty based on PFC activity before training. The model achieved 86% accuracy in predicting the appropriate difficulty level. After three weeks of training, participants showed improved executive function and reduced PFC activation, indicating increased neural efficiency (Jeun et al., 2022). However, the study was limited by a small sample size and lacked DeepNet approaches for improved accuracy.
Pan et al. applied ML techniques to classify instructional approaches based on brain-to-brain coupling between instructors and learners. They found that machine learning models could distinguish between scaffolding-based learning and explanation-based learning better than single-brain models that highlights the role of interactive engagement in learning efficiency (Pan et al., 2020).
Another approach was proposed by Saikia, where k-means clustering has been used to classify individuals using fNIRS data during a working memory task. Participants performed n-back tasks at four difficulty levels, and clustering revealed three distinct response groups. The study also introduced inverse efficiency score (IES) as a refined performance metric that provided clearer distinctions between difficulty levels compared to traditional methods (Saikia, 2023). However, the model faced challenges in generalizability due to individual differences in brain responses.
Oku and Sato used fNIRS data to model with random forests and logistic regression to classify student engagement and predict quiz performance in online learning platforms. The random forest model achieved a ROC-AUC of 0.67, while logistic regression achieved 0.65, identifying specific brain regions (F4-F6 and AF3-AFz) as key predictors of performance (Oku and Sato, 2021).
3. Experiment
3.1. User Study
Figure 2. Overview of the study procedures: Participants first received an introduction to the study and completed a demographic questionnaire. A training session was provided to familiarize them with the experimental interface and task format. The main session began with a 20-second resting baseline, followed by the first task segment consisting of four questions. After a short 20-second rest, the second segment commenced. This was followed by a longer break and the administration of post-task questionnaires such as the post-test and user experience (UX). The same procedure was repeated for a second session after 10-12 minutes break.
[…]
Figure 2: Overview of the study procedure. Overview of the experimental protocol. Participants first received an introduction to the study and completed a demographic questionnaire. A training session was provided to familiarize them with the experimental interface and task format. The main session began with a 20-second resting baseline, followed by the first task segment consisting of four questions. After a short 20-second rest, the second segment commenced. This was followed by a longer break and administration of post-task questionnaires. The same procedure was repeated for a second session after 10-12 minutes break.The entire protocol was designed to assess participant responses across structured task blocks interspersed with rest periods.
Figure 2 represents the steps of the user study. The participants used a unity-based educational game. The game was designed to evaluate participants’ understanding of the basic structure of graph concepts, such as nodes, edges, loops, etc. Throughout the study, we recorded the participants’ quiz scores and hemodynamic responses from the fNIRS. The Qualtrics survey platform was used to collect responses to demographic questionnaire and pre and post-test questions. The study was approved by the institutional review board.
3.1.1. Apparatus
Figure 3. (a) Experimental setup: Participant playing an educational game wearing fNIRS headband. This setup captures hemodynamic responses in the prefrontal cortex, with data acquisition managed through COBI software on a secondary laptop. (b) 16 channels used for collecting $\Delta HbO$
[…]
Figure 3: (a) Experimental setup: Participant playing an educational game on one laptop wearing fNIRS headband. The fNIRS device is attached to another computer, which captures hemodynamic responses in the prefrontal cortex, with data acquisition managed through COBI software on a secondary laptop. (b) We can see an image of a brain. and pointing 16 channel of pre-frontal cortex that we used for collecting $\Delta HbO$
Figure 3 presents study setup and fNIRS headband orientation. We used two computers and an fNIRS device (Figure 3 (a)). We used an 18-channel fNIRS imager 2000S (fNIRS Device LLC, Potomac, MD, USA), 4 light emitters, wavelength ($\lambda$) 730-850 nm with average power $<1$ mW) sampled at 10 Hz. The detectors were separated by 2.5 cm with a penetrating depth of approximately 1.2 cm. This headband was placed on participants’ foreheads, and fNIRS collected data for 16 channels (Figure 3(b)). The user played the educational quiz games on a Dell XPS touchscreen laptop (L1) [Intel (R) Core(TM) i7-7700HQ processor at 2.80 GHz]. Another desktop computer (L2) [Intel(R) Core(TM) i7-10700T processor at 2.00 GHz] was connected to the fNIRS device. The laptop was used to collect fNIRS data using Cognitive Optical Brain Imaging (COBI) Studio Software. We also used fNIRSoft Software (Version 4.9) to apply filters on $\Delta HbO$ data and generate .CSV file for further processing.
3.1.2. Study Procedure
After signing the consent form and filling out the demographic questionnaires, the participants first experienced a demo version where the study was clearly explained and instructions about sitting still as much as possible (see  Figure 2).
After appearing on a ten-question pre-test, they watched the recorded video tutorial where they learned the basic terminologies of the graph. In our study, we divided our 16 questions into two sessions, with each session having two segments and each segment having four questions. Participants had up to 30 seconds to answer each question, and after each question, they had five seconds of feedback. So, each segment was 140 seconds $((4\times 30=120+4\times 5=20)=140)$ long. In between the two segments, there was a 20-second rest period. After the first session, they had a 6-10 minute break. After each session, the participants completed a post-test on the same graph topic. We counterbalanced the session sequence of the questions to mitigate learning curves.
3.1.3. Cognitive Effort Measure
In this study, we aimed to estimate the RNE and RNI in each session segment. We determined the overall average and standard deviation (SD) for each session (segment 1 & 2 combined, segment 3 & 4 combined).
To calculate RNE and RNI, we took the average of $\Delta HbO$ to measure oxygenated hemoglobin concentrations in the blood during the learning period in the overall PFC.The $\Delta HbO$ and performance score values were converted to Z-scores to reduce individual variability and to bring effort and performance onto the same scale which makes comparisons and RNE/RNI calculations more interpretable. In this study, we used the mean score as performance ($P_{z}$) and the inverse mean of $\Delta HbO$ as cognitive effort ($CE_{z}$) because we want better performance with minimum cognitive effort (Paas et al., 2003; Shewokis et al., 2015; Getchell and Shewokis, 2023; Koiler et al., 2022; Sharmin et al., 2024a). The participants can answer all questions correctly. To avoid division by zero when the standard deviation of the performance score became zero (i.e., in cases of uniform performance), we added a small epsilon constant ($\epsilon=0.001$) to the denominator of the z-score equations. This ensured numerical stability and preserved correct directional trends in RNE and RNI calculations (Equation 4).
(1) $\hskip 56.9055ptP_{z}=\dfrac{Score_{i}-Score_{\text{(GM)}}}{Score_{\text{(SD)}}+\epsilon}\hskip 28.45274pt$
[…]
(3) $\hskip 56.9055ptRNE=\dfrac{P_{z}-CE_{z}}{\sqrt{2}}\hskip 28.45274pt$
(4) $\hskip 56.9055ptRNI=\dfrac{P_{z}+CE_{z}}{\sqrt{2}}\hskip 28.45274pt$
3.2. Data Acquisition
We collected fNIRS data using COBI, and the fNIRS recording rate was 10 samples per second. The data was saved in .fNIRS format. Each question was presented for 30 seconds, so each file contained 300 rows for each question corresponding to 16 optode ($\Delta HbO$) values. Sixteen participants answered 16 questions each, leading to 256 responses in the dataset ($16\times 16$). The dataset was imbalanced because 168 responses were correct (class 1) and 88 were incorrect (class 0). This imbalance occurred because participants generally performed well, which suggests that task difficulty may need future adjustment for broader classification analyses. The dataset has 1228800 data points before pre-processing ($participants(16)\times questions(16)\times fNIRSpoint(300)\times optodes(16)$). Since fNIRS captures $\Delta HbO$) changes over time, the data is temporal in nature.
3.3. Data Pre-processing
After data acquisition, we pre-process our data using fNIRSoft software (see Figure 4). We checked raw light intensities and individual optodes visually.
Next, we used a finite impulse response filter (20th order, Hamming window), a low pass filter, which removed the raw light intensity data at 0.1 Hz input from physiological signals, such as respiration and heartbeat. Then we applied the Modified-Beer Lambert rule, where fNIRS transforms light intensity measurements into changes in the concentrations of $\Delta HbO$ (Cakir et al., 2015; Çakır et al., 2016). According to the rule, a light’s attenuation depends on the medium’s characteristics it passes through (Cakir et al., 2015; Çakır et al., 2016). A portion of the near-infrared light used in fNIRS is absorbed by hemoglobin in the blood arteries in the brain. oxygenated hemoglobin ($\Delta HbO$) concentrations are connected to changes in brain activity. fNIRS can quantify how much light is absorbed by calculating changes in $\Delta HbO$ (Cakir et al., 2015; Çakır et al., 2016). Finally, the detrending filter was applied to data characterizing changes in concentration to remove drift in the data using linear parameters that convert the baseline slope to zero. We got oxygenated hemoglobin $\Delta HbO$ from the available biomarkers. We labeled the data using fNIRSoft software and saved them in CSV file format.
Figure 4. Pre-processing and artifact removal of fNIRS data.
[…]
Figure 4: This figure shows the pre-processing steps to remove noise and artifact removal. In the first box, we checked raw light intensities and individual optodes visually. In the next text box, we used a finite impulse response filter (20th order, Hamming window), a low pass filter, where the raw light intensity data at 0.1 Hz input from physiological signals, such as respiration and heartbeat. For the third box, we applied the Modified-Beer Lambert rule, where fNIRS transforms measurements of light intensity into changes in the concentrations of $\Delta HbO$ and $\Delta HbO$. In the fourth box, the detrending filter was applied to data characterizing changes in concentration to remove drift in the data using linear parameters that convert the slope of the baseline to zero. In the fifth box, we labeled the fNIRS data based on task and time. and in the last box, we extract the $\Delta HbO$ in a .CSV file
Algorithm  LABEL:algo:_pre-process shows the steps of pre-processing used to prepare the fNIRS dataset for the proposed DeepNet model. The pipeline includes data extraction, missing value handling, standardization, feature selection, class balancing, reshaping, and train-test splitting. These steps ensure the dataset is clean, well-structured, and optimized for GRU-based model training.
At first, we extracted the quiz responses based on filenames. Then, we handled missing values by imputing with the mean per question and cleaned column names to prevent errors. Since participants often answered within 20 seconds, only the first 200 rows per question were retained. We grouped the data for performance (participant and question) prediction. After applying standardization (Z-score normalization) to ensure uniform feature scaling, we used Principal Component Analysis (PCA) to reduce dimensionality from 16 to 12 features (see LABEL:fig:PCA) to improve efficiency.
Finally, the dataset was split into 80% training and 20% testing. We had 256 quiz responses in total. After splitting, we had 208 responses (13 participants) for the training set and 48 (3 participants) for the test set. We had 134 data for correct answers (class 1) and 74 for wrong answers (class 0) in the training set. To balance this dataset, we applied the Synthetic Minority Over-Sampling Technique (SMOTE), which balances class distribution by oversampling the minority class. The data was then reshaped into (samples, 1 timestamp, 12 features) to be compatible with the GRU model.
4. Model Architecture
Our proposed hybrid DeepNet model (CNN-GRU) takes a reshaped PCA-processed input of shape (1, 12) per sample. The pipeline includes: A Conv1D layer with 32 filters and a kernel size of 1 to extract spatial patterns, a GRU layer with tunable hidden units (8 or 16) to capture temporal dependencies, a BatchNormalization and Dropout to reduce overfitting, and a fully connected dense layer of 64 units followed by a softmax output layer for binary classification. The CNN identifies which brain regions are most active during different cognitive states represented by fNIRS optodes. This helps the model to detect patterns related to effort.The GRU layer handles temporal dependencies in hemodynamic signals over the 30-second question window. This makes it well-suited to capture cognitive effort changes over time. This hybrid architecture is thus ideal for fNIRS data, which is both spatially distributed and temporally dynamic.
Training Strategy: We used an Adam optimizer with learning rates ranging from 0.0005 to 0.003, batch sizes from 1 to 32, and early stopping based on validation loss with patience of 8 epochs. For each hyperparameter combination, the model was trained for a maximum of 150 epochs, with the best model checkpoint saved based on validation accuracy.
Post-Training Analysis: Following model training, GRU embeddeddings were used as input features to an XGBoost classifier to further evaluate learned representations. For each model variant, we report Train Accuracy, Test Accuracy, Precision, Recall, F1-Score, and XGBoost Accuracy on the extracted features.
Hyperparameter Search A grid search over 72 configurations was conducted, varying:
GRU units: 8, 16; Dropout rates: 0.1, 0.2, 0.4; Learning rates: 0.0005, 0.001, 0.003; Batch sizes: 1, 4, 8, 16, 32.
All results were logged, and models were evaluated on an unseen test set. We selected the hyperparameter based on validation accuracy to ensure optimal performance.
Figure 5. DeepNet Model Architecture. We fed fNIRS data into our model and to get performance score as output
[…]
Figure 6: DeepNet Model Architecture. The flow shows the dataset goes through the CNN layer, batch normalization, GRU layer, dropout(0.1), then dense layer, and finally the Softmax layer to get the final output.
This hybrid architecture captures spatial, temporal, and contextual relationships in the features extracted from fNIRS signals (see Table 1.
Table 1. Hyperparameters Used in the Model
[…]
Table 1: This table shows the Hyperparameter Used in the Model. It has two column parameters and its value. Hyperparameter Value Learning Rate 0.003 GRU Units 8 Convolutional Filters 32 Dropout Rate 0.1 Batch Size 4 Optimizer Adam Loss Function Categorical Cross-Entropy
Performance Metrics To evaluate our model, we compute precision, recall, and F1-score. Precision measures the proportion of correctly predicted positive observations to the total predicted positive observations. Recall evaluates the proportion of correctly predicted positive observations to all actual positive observations. A high recall indicates the model successfully captures most of the positive instances. F1-score is the harmonic mean of precision and recall. It is particularly useful when the class distribution is imbalanced.
5. Results
In the following sections, we explain how different models performed to detect performance scores and how these predicted score works to estimate cognitive effort.
5.1. Prediction of Performance Score
At first, we ran our dataset through traditional ML algorithms. Table 2 presents the evaluation metrics for different ML models based on 5-fold cross-validation. In most cases, we obtained higher training accuracy than test accuracy. The Random Forest model achieved the highest recall ($0.92$) and F1-score ($0.77$). SVM and XGBoost also demonstrated strong recall values ($0.82$ and $0.80$, respectively), with SVM achieving an F1-score of $0.73$. Gradient Boosting and CatBoost showed balanced precision-recall values, which indicates stable performance across categories. However, Neural Network and LightGBM showed poor generalization with the lowest test accuracy values ($0.65$ and $0.54$, respectively). Overall, models like Random Forest, SVM, and XGBoost exhibited higher sensitivity and robustness, while Neural Network and LightGBM struggled with consistency in classification. Also, the models were overfitted.
After evaluating traditional methods, we applied the dataset to various deep-learning models. Table 2 (bottom section) shows the performance comparison of different DeepNet architectures for the binary classification task. Our proposed hybrid model, CNN-GRU, outperformed all baseline models. It achieved a test accuracy of 73.08%, with a remarkable F1 score of 82.05%, demonstrating excellent performance in balancing precision (73.81%) and recall (91.18%). The notable improvement in recall indicates that the CNN-GRU model is particularly effective at minimizing false negatives.
Among the baseline DL models, BiLSTM showed moderate performance with a test accuracy of 66.18% and an F1 score of 66.11%. After training the deep models, we fed the extracted features into an XGBoost classifier. While XGBoost on raw features gave reasonable performance, its accuracy improved when applied to deep-learned representations. When we trained an XGBoost classifier using features extracted from these deep models, BiLSTM and CNN achieved an XGBoost accuracy of 66.18%, while LSTM features gave a lower accuracy of 58.82%. The highest XGBoost accuracy (69.23%) was obtained using features from the hybrid CNN-GRU model. This result highlights the usefulness of deep features and shows that combining CNN (for spatial feature extraction) and GRU (for temporal modeling) enhances representation learning—even for models like XGBoost.
Table 2. Performance Comparison of Machine Learning and DeepNet Models
[…]
Table 2: compares the performance of machine and DeepNet models. the result shows that the proposed CNN-GRU model gives better accuracy to predict performance score. Model Train Test Precision Recall F1 Score ML Models Random Forest 1.00 0.63 0.66 0.92 0.77 Gradient Boosting 1.00 0.59 0.60 0.74 0.70 AdaBoost 0.98 0.57 0.66 0.73 0.69 SVM 0.70 0.64 0.65 0.82 0.73 Neural Network 0.71 0.65 0.63 0.70 0.66 LightGBM 1.00 0.54 0.59 0.73 0.68 CatBoost 1.00 0.64 0.65 0.81 0.72 DL Models CNN 0.68 0.56 0.56 0.56 0.56 LSTM 0.68 0.60 0.61 0.60 0.60 BiLSTM 0.72 0.66 0.66 0.66 0.66 CNN-GRU* 0.78 0.73 0.74 0.91 0.82
We conducted paired t-tests comparing CNN-GRU with each model using 5-fold test accuracies. While CNN-GRU performed better than others, none of the differences were statistically significant ($p>0.05$). Yet, these findings indicate that DeepNet models like the proposed CNN-GRU hybrid offer better generalization and classification capability for this time series task. This also suggests that learning directly from data using deep models provides better predictive performance compared to traditional approaches.
5.2. Feature Attribution Analysis (Interpretability)
To investigate which brain regions most influenced the model’s predictions, we conducted a multi-stage interpretability analysis combining PCA, GRU feature correlation, and SHAP explanations. 1. PCA-Optode and Region Mapping: we aggregated optode contributions by anatomical region LPFC and VMPFC to obtain regional level insight. PC2 exhibited a stronger total contribution from LPFC $(6.99)$ compared to VMPFC $(5.36)$, which suggests that this component reflects distributed LPFC activity.
2. GRU - PCA Correlation (Feature Encoding Stage) We then analyzed how GRU latent features (from the encoding stage) relate to the PCA components. As shown in Figure 6, GRU Feature 6 had the highest correlation with PC2 $(r=0.65)$. Given PC2 was previously found to be LPFC-dominated, this suggests GRU6 is encoding task-relevant signals originating from LPFC.
Figure 6. Correlation between GRU features and PCA components
3. SHAP - GRU Feature Importance (Decision Stage) Finally, we applied SHAP (SHapley Additive exPlanations) to identify which GRU features most influenced the model’s final predictions. As shown in Figure 7, GRU Features 3 and 0 had the highest SHAP values, indicating their strong impact in decision-making. Interestingly, these features also showed notable correlation with PC2—GRU3 negatively $(r=-0.67)$ and GRU0 positively $(r=0.42)$.
Figure 7. SHAP summary plot showing GRU feature importance in prediction
5.3. Estimation of Cognitive Effort
After identifying the proposed CNN-GRU as the best-performing model, we used its predicted performance scores to calculate RNE and RNI. These metrics provided a deeper understanding of participants’ cognitive states across task segments. Our test set included three participants: P8, P11, and P16.
Table 3 presents a detailed comparison of actual and predicted RNE and RNI values across task segments for three test participants. We calculated the RNE and RNI from the predicted score from the deep learning model ( Equation 4). The final row shows statistical similarity using Mean Absolute Error (MAE) and Pearson correlation. The very low MAE and near-perfect correlation $(r=0.999)$ confirm that the model’s predicted RNE and RNI values closely follow the actual measurements.
(a) RNE Comparison
(b) RNI Comparison
Figure 8. Line plots of predicted and actual RNE and RNI values across task segments for each test case.
[…]
Figure 7: Line plots of predicted and actual RNE and RNI values across task segments for each test case. The lines show that the predicted line follows the actual one.
Table 3. Comparison of Actual and Predicted RNE and RNI across Participants and Task Segments. Predicted values closely follow the actual values, as confirmed by low MAE and high Pearson correlation.
[…]
Table shows actual and predicted values of RNE and RNI for participants P8, P11, and P16 across 4 segments. A summary row shows statistical similarity using Mean Absolute Error (MAE) and Pearson correlation coefficient. Participant Segment RNE RNI Actual Predicted Actual Predicted P8 1 0.14 0.9 -0.63 0.13 2 0.25 -0.51 0.25 -0.51 3 -32.65 -32.98 32.65 32.32 4 -0.0002 0.33 0.0002 0.33 P11 1 0.95 0.53 -0.61 -1.03 2 -0.17 0.25 -0.17 0.25 3 6.80 6.94 -7.57 -7.44 4 0.38 0.25 0.38 0.24 P16 1 -2.33 -2.46 -1.83 -1.70 2 0.25 0.38 -0.25 0.38 3 -0.59 -0.59 -0.59 -0.59 4 0.0002 0.0002 0.0002 0.0002 Mean absolute error (MAE) 0.29 0.38 Pearson Correlation 0.99 0.99
To understand participants efficiency and involvement we plot $CE$ in X axis and $P_{z}$ in Y axis (see  Equation 4). Figure LABEL:fig:RNE_RNI_ML_dynamic visualizes participant-wise relative neural efficiency (RNE) and involvement (RNI) in Cartesian quadrants representing four cognitive states: HE+HI (High Effort, High Involvement), HE+LI (High Effort, Low Involvement), LE+HI (Low Effort, High Involvement), and LE+LI (Low Effort, Low Involvement). Predicted coordinates from the DeepNet model were plotted alongside the actual values for three participants (P8, P11, P16) across four task segments.
Out of the 12 plotted points, we can see that 09 predicted coordinates matched the actual cognitive states. This highlights the accuracy and interpretability of our CNN-GRU-based performance prediction model for estimating RNE and RNI.
In segment 1, participants began in LE+HI or LE+LI regions, which suggests that their initial performance was high with low cognitive effort, which is considered a potentially favorable cognitive state for learning. In segment 2, while performance levels varied across participants, most points clustered around the origin, which indicates moderate cognitive effort and performance, and suggests a transitional cognitive state.
In segment 3, we noticed a drop in cognitive effort for P11 and P16. This could be the reason for the longer break after segment 2. P8 reported having a headache after segment 2, which could be the reason for increased effort in segment 3. However, in segment 4, the effort dropped after a short rest for P8. In the other two cases, in segment 4, participants maintained a general transition of cognitive state.
6. Discussion
The aim of this study is to estimate cognitive effort during educational tasks by analyzing fNIRS-derived hemodynamic responses and predicted performance scores. We proposed a hybrid CNN-GRU DeepNet model to predict task performance from brain signals and then we derived Relative Neural Efficiency (RNE) and Relative Neural Involvement (RNI) to interpret learner efficiency and involvement. The contribution of this work is that it shows that cognitive effort can be meaningfully measured using a combination of brain signals and deep learning-based performance prediction. These findings might influence future adaptive learning systems.
For RQ1, the results suggest that the DeepNet model performed better than traditional machine learning models. It also shows that using CNN, LSTM, and BiLSTM alone does not give the best results for score prediction using fNIRS. The hybrid CNN-GRU model achieved the highest accuracy (73.08%). This suggests that a combination of CNN for feature extraction and GRU helps improve classification performance.
The inclusion of fNIRS signals with DeepNet enabled the estimation of cognitive metrics such as RNE and RNI from predicted performance scores. For RQ2 these derived measures help interpret learners’ cognitive effort, involvement, and efficiency during educational gameplay.
These findings highlight the importance of personalized learning design. We believe the behavior of RNE and RNI is influenced by several cognitive and physiological factors. For example, CE dropped down after a post-break, which aligns with the hypothesis that rest helps restore executive functioning. Also, as participants became more familiar with the task structure, learning progression probably helped to improve efficiency and involvement. These transitions suggest that both task familiarity and rest contribute to variations in neural efficiency and involvement over time.
For RQ3 these derived measures help interpret learners’ cognitive effort, involvement, and efficiency during educational gameplay.By visualizing predicted and actual scores alongside cognitive effort, instructors and researchers can adjust instructional strategies and interventions more precisely. Results suggest that RNE and RNI are not static but respond to breaks, fatigue, and learning progression. Breaks consistently led to performance recovery, as seen in increased RNE post-break. For RQ3 these derived measures help interpret learners’ cognitive effort, involvement, and efficiency during educational gameplay. Even with moderate accuracy, the predicted CE trends followed actual RNE/RNI patterns. Since we use aggregated and z-scored values for CE estimation, occasional false positives/negatives have a limited impact on the overall cognitive state trends. The derived RNE/RNI values smooth out individual prediction errors. Also, to observe cognitive effort patterns, we used a Cartesian plot. We observed the predicted values most of the time aligned with the actual values in the same quadrant of the Cartesian plot. This spatial overlap indicates that the model was able to capture similar patterns of cognitive effort behavior. Therefore, the proposed model not only supports cognitive effort estimation but also aids in the interpretation of cognitive state trends across different learning segments.
However, future improvements can reduce this noise further. Our goal was not to maximize classification accuracy alone, but to enable interpretable CE estimation using RNE and RNI. The CNN-GRU model offered a good tradeoff between accuracy and signal pattern learning. We plan to explore more robust or adaptive architectures in future work. CE is highly individualized. While our model captures general trends, it does not yet account for individual baselines or variability in fNIRS responses. Adaptive or personalized calibration may improve robustness in future work.
Limitations and Future Work
Real-time implementation was beyond the scope of this study, but the CNN-GRU model we used is lightweight. So, it should be able to make predictions in real-time. In the future, deploying this system in real-time during learning sessions could help detect fatigue and adjust task difficulty dynamically, contributing to more effective and personalized educational experiences. To improve generalizability, future work will explore data augmentation (e.g., jittering, time warping of signals) and transfer learning from larger fNIRS datasets. We also plan to extend the study to diverse populations in classroom environments to validate the model. Additionally, adjusting the duration of learning tasks, such as extending or shortening session lengths, may provide further insight into how cognitive effort and motivation evolve. Exploring the impact of different types of learning materials (e.g., visual vs. interactive, gamified vs. traditional) on RNE and RNI could also uncover new instructional design principles.
7. Conclusion
This study demonstrates a method for estimating and interpreting cognitive effort through performance scores predicted by DeepNet and oxygenated hemoglobin ($\Delta HbO$) in the brain collected using fNIRS. In this work, we proposed a hybrid model to predict performance scores in an educational game. We collected $\Delta HbO$ using fNIRS and fed it into our CNN-GRU model for the prediction. This predicted score was used to estimate cognitive effort by measuring RNE and RNI. Our hybrid model showed 73.08% accuracy with 82% F1-Score. After estimating RNE and RNI, we observed that the predicted RNE and RNI closely follow the actual RNE and RNI. This approach paves the way for a personalized educational setup that can dynamically adapt to learners’ cognitive states, potentially transforming classroom and remote educational settings. By deriving RNE and RNI from predicted performance and effort, we reveal meaningful trends tied to cognitive states. The findings highlight the model’s potential in adaptive learning contexts and highlight future directions in real-time neuro-adaptive systems.
Acknowledgments
We wish to express our gratitude to the study participants and lab members.